OpenStudy (anonymous):

what is the 6th term in 240,120,60

OpenStudy (anonymous):

use arithmetic progression

hartnn (hartnn):

hi, do you know about geometric progression ?

OpenStudy (anonymous):

Using arithmetic progression i got: -600 With Geometric: 7.5 Arithmetic: A(n)=A(1) + (n-1)d = 240 + (5*-.5) Geometric A(n)=A(1)*r^(n-1) = 240*(.5)^(5)

hartnn (hartnn):

its not in Ap, its in GP :)

OpenStudy (anonymous):

I personally think it will be Geometric.

OpenStudy (anonymous):

Because of the ratio right?

hartnn (hartnn):

yes, we have a common ratio between the terms.

OpenStudy (31356):

Go to: {example} http://answers.yahoo.com/question/index?qid=20110804090812AA7HSHT

OpenStudy (31356):

Medal?

OpenStudy (anonymous):

Why are you asking for a medal for googling something for him/her?

OpenStudy (♪chibiterasu):

If he would've gotten the answer from Yahoo Answers then he may as well be in Yahoo and no here. .-.

OpenStudy (anonymous):

lmfao